A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. n = 40, p = 0.02, x-2 P(2)
Answer #1

Solution

Given that ,

p = 0.02

1 - p = 1 - 0.02 = 0.98

n = 40

x = 2

Using binomial probability formula ,

P(X = x) = ((n! / x! (n - x)!) * px * (1 - p)n - x

P(X = 2) = ((40! / 2! (38)!) * 0.022 * (0.98)38

= 0.1448

P(2) = 0.1448